# Demystifying Spectral Feature Learning for Instrumental Variable Regression

### v2

@@ 5 Experiments
The previous experiment assumes $h_{0}$ lies in the span of the singular functions of $\mathcal{T}$. Figure˜2 illustrates three regimes. The left panel shows a well-aligned case (“good/bad”), where $h_{0}$ is entirely supported on the singular functions of $\mathcal{T}$, allowing for exact recovery as long as enough data is available. The middle panel represents a “partial alignment” regime, where $h_{0}$ has some overlap with the singular functions; in this case, recovery is imperfect, but the estimator can still capture some signal. Finally, the right panel illustrates the “misaligned” or “very ugly” regime, where $h_{0}$ is orthogonal to the singular functions. The instrument provides no information about the signal, and all spectral methods fail. We refer to Appendix C for further discussion of these scenarios and the experimental setup.
Appendix A Proofs
@@ A.1 Proofs of Section˜2
In order to prove ˜1, we use the following lemma.
@@
We recall and prove ˜1. This is a classical result, see, e.g., [25].
See 1
@@
$=\sum_{i\geq 1}\int_{\mathcal{Z}}\langle e_{i},p(\cdot,z)\rangle_{L_{2}(X)}^{2}\,d\pi_{Z}(z)$ (by Eq.˜13)
@@ A.2 Proofs of Section˜3
˜4 is due to Lemma 1 of [3] and ˜2 corresponds to part (iii). We provide a full proof for completeness.
@@
We recall and prove Corollary˜1:
See 1
@@
This implies, using that $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by ˜2,
@@
Since both parts of Assumption˜3 hold, and $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by ˜2, we may apply Theorem˜1, which yields the desired result. ∎
A.3 Proofs of Section˜4
@@
$\Rightarrow$ Assume $p(x,z)=\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)$. Then, by Eq.˜10, for all $h\in L_{2}(X)$,
@@
In order to prove Theorem˜3 we prove intermediate results.
@@
We first show that $\sigma_{d}^{-1}\leq\tau_{\hat{\varphi},d}\leq(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}$. First note that if $h\in\mathcal{H}_{\hat{\varphi},d}$, then by Proposition 7, $h\in\mathcal{V}_{d}$. Using the SVD of $\hat{\mathcal{T}}_{d}$ from Eq.˜14, and writing $h=\sum_{i=1}^{d}\langle h,\hat{v}_{i}\rangle\hat{v}_{i}$, we obtain:
@@
We now prove the equivalence between Eq.˜6 and Eq.˜7. While the proof strategy follows that of Kostic et al. [24, Theorem 1], we include our own version for completeness, as the parametrization of the learned operator differs. Specifically, [24] directly approximate the truncated operator $\mathcal{T}_{d}$ using a singular value decomposition of the form $\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=2}^{d}\hat{\sigma}_{i}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$, where the singular values $\hat{\sigma}_{i}$ are learned explicitly via a separate network. In contrast, our approach learns only the feature maps $(\varphi_{i},\psi_{i})$ and defines the approximation $\mathcal{T}_{d}(\hat{\varphi},\hat{\psi})=\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$. As such, we provide a short self-contained derivation adapted to our setting for clarity. Recall that the Eckart–Young–Mirsky formulation of the objective is defined as
@@
Appendix B Extended Related Work
@@
Building on similar principles, [17] proposed nonparametric methods based on both kernel444”Kernel” is meant here in the sense of density estimation, and not reproducing kernel: see [17, eq. (2.4)]. techniques and orthogonal series for estimating regression functions with instrumental variables. For their orthogonal series estimator, they transformed the marginal distributions of $X$ and $Z$ to be uniform and used Fourier expansions. The estimated Fourier coefficients of the structural function were obtained via a regularized regression involving estimated coefficients from the first stage. They were among the first to derive optimal convergence rates for this class of problems, explicitly linking these rates to the “difficulty” of the ill-posed inverse problem, which is characterized by the eigenvalues of the underlying integral operator.
@@
[33] introduced Kernel Instrumental Variable Regression (KIV), a direct nonparametric generalization of 2SLS. KIV models the relationships between the different variables as nonlinear functions in reproducing kernel Hilbert spaces (RKHSs). In stage 1, it learns a conditional mean embedding $\mu(z)=\mathbb{E}\left[\varphi(X)\mathrel{}\middle|\mathrel{}Z=z\right]$ where $\varphi(X)$ represents features of $X$ mapped into an RKHS $\mathcal{H}_{X}$. This learning is framed as a vector-valued kernel ridge regression, effectively estimating the conditional expectation operator $E:\mathcal{H}_{X}\to\mathcal{H}_{Z}$. In stage 2, a scalar-valued kernel ridge regression of the outcome $Y$ on the estimated means $\hat{\mu}(Z)$ is performed to estimate the structural function. The authors prove the consistency of KIV in the projected norm under mild conditions, and derive conditions under which KIV achieves minimax optimal convergence rates. The analysis was later improved by [29], who establish convergence in $L_{2}$ norm rather than the projected norm.
B.0.0.0.2 Saddle-point approaches.
@@
B.0.0.0.3 About the contrastive loss.
@@
Appendix C Experimental Details
In this section, we provide additional details for the experiments presented in the main text. Let $(Z,X)$ be generated as described in Section˜5 with $n=100000$. Figure˜3 displays both the true data-generating density and the density corresponding to a set of learned spectral features. Even for $d=11$, the resulting densities are complex. While it is feasible to conduct analogous experiments with higher values of $d$, we observed no qualitative changes in the outcomes. However, training and hyperparameter tuning became increasingly challenging as $d$ grew.
@@ C.1 Models Employed
C.2 Expanded example of the ugly scenario
@@
mirroring the setup from Section˜5. We vary the spectrum by setting $\sigma_{1:k}=c$ and $\sigma_{k+1:d-1}=0$ for $k\in\{1,\ldots,d-1\}$. We then define the target function:
@@
which is uniformly spread across the feature directions. As $k$ increases, the projection of $h_{0}$ onto the singular space of $\mathcal{T}$ increases in discrete steps from 0 to 1. This results in a corresponding qualitative improvement in the accuracy of the 2SLS estimator for $h_{0}$. Figure˜4 illustrates this behavior: for small values of $k$, where $h_{0}$ is largely orthogonal to the singular functions of $\mathcal{T}$, the estimate fails to recover $h_{0}$. As $k$ grows, and more of $h_{0}$’s energy lies in the span of these singular functions, the 2SLS estimate increasingly aligns with the true function.
@@
Appendix D Estimating spectral alignment
In this section, we present a methodology to estimate the spectral decay of the operator $\mathcal{T}$ as well as spectral alignment with the structural function $h_{0}$. Recall that $\mathcal{T}$ is compact and admits an SVD of the form
$\mathcal{T}=\sum_{i\geq 1}\sigma_{i}u_{i}\otimes v_{i},$
with the key relationship that $\sigma_{i}v_{i}=\mathcal{T}^{*}u_{i}$. Therefore, for $i=1,\ldots,d$
$\mathbb{E}[v_{i}(X)h_{0}(X)]=\langle v_{i},h_{0}\rangle_{L_{2}(X)}=\frac{1}{\sigma_{i}}\langle\mathcal{T}^{*}u_{i},h_{0}\rangle_{L_{2}(X)}=\frac{1}{\sigma_{i}}\langle u_{i},r_{0}\rangle_{L_{2}(Z)}=\frac{1}{\sigma_{i}}\mathbb{E}[Y\cdot u_{i}(Z)].$ (15)
To evaluate spectral alignment, we shall approximate the RHS of the above. Suppose we have learned an operator $\widehat{\mathcal{T}}_{d}=\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}=\Psi\Phi^{*}$ which approximates the true conditional mean operator $\mathcal{T}$. Here $\Psi\colon\mathbb{R}^{d}\to L^{2}(Z)$ and $\Phi\colon\mathbb{R}^{d}\to L^{2}(X)$ are the maps sending $\alpha\stackrel{{\scriptstyle\Psi}}{{\mapsto}}\sum_{i=1}^{d}\alpha_{i}\hat{\psi}_{i}$ and $\alpha\stackrel{{\scriptstyle\Phi}}{{\mapsto}}\sum_{i=1}^{d}\alpha_{i}\hat{\varphi}_{i}$. One can compute the SVD of $\widehat{\mathcal{T}}_{d}$ in two steps. We start with the derivations assuming access to population quantities. Let $C_{Z}$, and $C_{X}$ be the $d\times d$ population covariance matrices of the learned $Z$, and $X$ features, respectively. For any $z$ and $x$, we then introduce the whitened feature vectors $\psi_{i}^{\prime}(z)=(C_{Z}^{-1/2}\hat{\psi}(z))_{i}$ and $\varphi_{i}^{\prime}(x)=(C_{X}^{-1/2}\hat{\varphi}(x))_{i}$. The corresponding operators, $\Psi^{\prime}=\Psi C_{Z}^{-1/2}$ and $\Phi^{\prime}=\Phi C_{X}^{-1/2}$, are isometries. We can then write

### v3

@@ 6.1 Synthetic datasets
The previous experiment assumes $h_{0}$ lies in the span of the singular functions of $\mathcal{T}$. Figure˜2 illustrates three regimes. The left panel shows a well-aligned case (“good/bad”), where $h_{0}$ is entirely supported on the singular functions of $\mathcal{T}$, allowing for exact recovery as long as enough data is available. The middle panel represents a “partial alignment” regime, where $h_{0}$ has some overlap with the singular functions; in this case, recovery is imperfect, but the estimator can still capture some signal. Finally, the right panel illustrates the “misaligned” or “very ugly” regime, where $h_{0}$ is orthogonal to the singular functions. The instrument provides no information about the signal, and all spectral methods fail. We refer to Appendix E for further discussion of these scenarios and the experimental setup.
6.2 Dsprites dataset
Figure 3: Example image from the dSprites dataset.
The dSprites dataset, as employed in the evaluation of IV models, consists of $64\times 64$ noisy images of hearts, with varying $x,y$ positions, orientations and sizes. The full setting is as follows; $X\in\mathbb{R}^{64\times 64}$ is the raw image, $Z\in\mathbb{R}^{3}$ consists of the $x$ position, orientation and scale of the heart in the image, and the output is constructed as [^4] The original formulation of the dSprites model for IV used $A_{i,j}\sim\mathcal{U}(0,1)$ [40]. However, it is noted in [41] (arxiv version 18/06/2024, Appendix E.3) that it leads to identifiability issues and that the choice $A_{ij}=|32-j|/32$ is preferred.
$Y=h_{0}(X)+32\cdot(y\text{ position}-32)+U,\,U\sim\mathcal{N}(0,1),$
where
$h_{0}(X)=(\|A\circ X\|^{2}-3000)/500,\,A_{ij}=|32-j|/32.$
It was observed by Sun et al. [36] that spectral methods significantly outperform alternatives in this setting. We argue that it is because the structural function employed in this benchmark lies in the “good” regime. Utilising the alignment estimation methods outlined in Section˜5, we are able to provide evidence for this claim. We observe that the structural function is spanned by the leading singular functions of the conditional expectation operator associated with this model. The details of this experiment can be found in Appendix˜F.
Acknowledgements
D.M., J.W., and A.G. are supported by the Gatsby Charitable Foundation. V.R.K. is supported from NextGenerationEU and MUR PNRR project PE0000013 CUP J53C22003010006 “Future Artificial Intelligence Research (FAIR)”. A.M. has received funding from the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 950180), and travel support from ELSA (Project ID: 101070617).
Supplementary Material: Demystifying Spectral Feature Learning for Instrumental Variable Regression
Appendix A Background on Spectral Theory
For completeness, we briefly recall key notions from spectral theory used throughout the paper. Our goal is to provide intuition and the minimal mathematical tools required to interpret concepts such as compactness, singular value decomposition (SVD) of operators, and eigensubspaces.
Let $H_{1}$ and $H_{2}$ be separable Hilbert spaces with inner products $\langle\cdot,\cdot\rangle_{H_{i}}$ and norms $\|\cdot\|_{H_{i}}$, $i=1,2$.
Hilbert spaces and bounded operators.
A linear map $A:H_{1}\to H_{2}$ is said to be bounded if there exists $C>0$ such that $\|Ah\|_{H_{2}}\leq C\|h\|_{H_{1}}$ for all $h\in H_{1}$. The smallest such constant is the operator norm, denoted $\left\|A\right\|_{\mathrm{op}}$.
Adjoint and self-adjoint operators.
Every bounded operator $A:H_{1}\to H_{2}$ admits an adjoint $A^{\ast}:H_{2}\to H_{1}$ defined by $\langle Ah,g\rangle_{H_{2}}=\langle h,A^{\ast}g\rangle_{H_{1}}$. If $H_{1}=H_{2}$ and $A=A^{\ast}$, the operator is self-adjoint.
Compact operators.
A linear operator $A$ is called compact if it maps bounded sets in $H_{1}$ to relatively compact sets in $H_{2}$, that is, the image of any bounded sequence has a convergent subsequence in $H_{2}$. All compact operators are bounded, but not all bounded operators are compact. Any compact operator is a limit (in operator norm) of finite-rank operators, so that the class of compact operators can be defined alternatively as the closure of the set of finite-rank operators in the operator norm topology. The conditional expectation operator $\mathcal{T}$ is a canonical example of a compact operator (see Assumption˜2 below).
Hilbert–Schmidt operators and norm.
Let $A:H_{1}\to H_{2}$ be a bounded linear operator. Choose any orthonormal basis $\{e_{j}\}_{j\geq 1}$ of $H_{1}$. The operator $A$ is called Hilbert–Schmidt if
$\left\|A\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}\coloneqq\sum_{j\geq 1}\|Ae_{j}\|_{H_{2}}^{2}<\infty.$
The value of the sum above is independent of the chosen basis of $H_{1}$. Hence the quantity $\langle A,B\rangle_{\scriptscriptstyle\mathrm{HS}}\coloneqq\sum_{j\geq 1}\langle Ae_{j},Be_{j}\rangle_{H_{2}}$ defines an inner product on the space of Hilbert–Schmidt operators, and $\|\cdot\|_{\scriptscriptstyle\mathrm{HS}}$ is the induced norm.
Singular Value Decomposition (SVD).
When $A:H_{1}\to H_{2}$ is compact (not necessarily self-adjoint), it admits a singular value decomposition:
$A=\sum_{i\geq 1}\sigma_{i}\,u_{i}\otimes v_{i},$
where $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq 0$ are the singular values of $A$ such that $\sigma_{i}\to 0$, and $\{v_{i}\}\subset H_{1}$, $\{u_{i}\}\subset H_{2}$ form orthonormal bases satisfying $Av_{i}=\sigma_{i}u_{i}$ and $A^{\ast}u_{i}=\sigma_{i}v_{i}$. This generalizes the matrix SVD to infinite-dimensional spaces. Important properties we use below:
  • The sum above is finite if and only if $A$ has finite rank.
  • $\left\|A\right\|_{\mathrm{op}}=\sigma_{1}$.
  • If $A$ is Hilbert–Schmidt then $A$ is compact such that $\left\|A\right\|_{\scriptscriptstyle\mathrm{HS}}^{2}=\sum_{i}\sigma_{i}^{2}<\infty$, hence $\left\|A\right\|_{\mathrm{op}}\leq\left\|A\right\|_{\scriptscriptstyle\mathrm{HS}}$. Thus Hilbert–Schmidt operators are bounded, but the converse need not hold.
Appendix B Learning Rate with Spectral Features
In this section, we formalize the learning rate that can be obtained from Definition˜2. To do so, we employ two standard assumptions from the nonparametric IV literature [8]: a source condition and a singular value decay assumption on the operator $\mathcal{T}$.
Assumption 6 (Singular Value Decay).
Let $p\geq 1$. There are constants $0<c_{1}\leq c_{2}<\infty$ such that for all $i\geq 1$, $c_{1}i^{-p}\leq\sigma_{i}\leq c_{2}i^{-p}$.
This assumption is standard in the nonparametric IV literature [8] and characterizes the difficulty of the inverse problem defined by the operator $\mathcal{T}$. A larger $p$ indicates faster decay and thus greater ill-posedness. Alternative ill-posedness characterizations can be considered, such as exponential decay [8].
Assumption 7 (Source Condition).
Let $r>0$. There is a constant $B<\infty$ such that
$\|h_{0}\|_{r}^{2}:=\big\|(\mathcal{T}^{*}\mathcal{T})^{-r/2}h_{0}\big\|_{L^{2}(X)}^{2}=\sum_{i\geq 1}\frac{\langle h_{0},u_{i}\rangle_{L^{2}(X)}^{2}}{\sigma_{i}^{2r}}\leq B.$
This condition quantifies how “smooth” $h_{0}$ is relative to the spectral decomposition of $\mathcal{T}$, or equivalently, how quickly the alignment coefficients $\langle h_{0},u_{i}\rangle_{L^{2}(X)}$ decay relative to the singular values. A larger $r$ implies $h_{0}$ is “smoother” or better aligned with the less ill-posed directions of $\mathcal{T}$. Note that under the “bad” scenario where the eigen-decay is fast ($p$ is large), a stronger alignment is necessary for $\|h_{0}\|_{r}^{2}$ to be finite for a large value of $r$. While in the “good” scenario with $p\sim 1$, $\|h_{0}\|_{r}^{2}<+\infty$ is more likely to hold with a large value of $r$.
Under Assumption 7, we have
$\|\left(\Pi_{\mathcal{X},d}\right)_{\perp}h_{0}\|_{L_{2}(X)}^{2}=\sum_{i>d}\langle h_{0},u_{i}\rangle_{L^{2}(X)}^{2}\leq\sigma_{d}^{2r}\sum_{i>d}\frac{|\langle h_{0},u_{i}\rangle_{L^{2}(X)}|^{2}}{\sigma_{i}^{2r}}\leq\sigma_{d}^{2r}\|h_{0}\|_{r}^{2}.$
Combined with Assumption 6, we obtain from Definition˜2 that:
$\|\hat{h}-h_{0}\|_{L^{2}(X)}=\mathcal{O}_{p}\left(\|h_{0}\|_{r}d^{-rp}+\sqrt{\frac{d^{1+2p}}{n}}\right).$
Balancing both terms in $d$ as a function of $n$, we obtain
$\|\hat{h}-h_{0}\|_{L^{2}(X)}=\mathcal{O}_{p}\left(n^{-\frac{r}{2(r+1)+p^{-1}}}\right),$
with
$d(n)=n^{\frac{1}{2p(r+1)+1}}.$
This bound is tight as a matching lower bound is obtained in Theorem 1-(i) [8]. From this tight rate, we see that both the ill-posedness ($p$) and the relative smoothness ($r$) fundamentally capture the effectiveness of 2SLS with spectral features, and their impacts are inherently intertwined.
  • A low relative smoothness (small $r$) means $h_{0}$ has significant components aligned with directions corresponding to very small singular values, even after accounting for the singular value decay. This also directly slows the rate, effectively rendering the problem very hard or even impossible ($r\to 0_{+}$ makes the rate vacuous, representing our “ugly” scenario).
  • A fast decay (large $p$) directly slows down the rate regardless of $r$. It means singular values drop quickly, making it fundamentally harder to resolve higher-frequency components of $h_{0}$. When $p$ is large we are either in the bad or in the ugly scenario. A large $p$ also slows down the rate indirectly by making it harder to achieve a high relative smoothness $r$.
When working with learned features.
We see from Assumption˜5 that when we learn the spectral features using contrastive learning, we incur an additional penalty related to $\hat{\varepsilon}_{d}/\sigma_{d}$ in the upper bound. This error shows that a faster eigendecay makes the 2SLS estimator more sensitive to errors in the learned features. Alignment of $h_{0}$ does not enter this factor, as learning the eigensubspaces of $\mathcal{T}$ is unrelated to $h_{0}$. Intuitively, features corresponding to small singular values can be empirically difficult to learn, even with a very large number of $(Z,X)$ samples.
Appendix C Proofs
@@ C.1 Proofs of Section˜2
In order to prove Assumption˜2, we use the following lemma.
@@
We recall and prove Assumption˜2. This is a classical result, see, e.g., [26].
[…]
@@
$=\sum_{i\geq 1}\int_{\mathcal{Z}}\langle e_{i},p(\cdot,z)\rangle_{L_{2}(X)}^{2}\,d\pi_{Z}(z)$ (by Eq.˜14)
@@ C.2 Proofs of Section˜3
˜3 is due to Lemma 1 of [3] and ˜1 corresponds to part (iii). We provide a full proof for completeness.
@@
We recall and prove Definition˜2:
[…]
@@
This implies, using that $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by ˜1,
@@
Since both parts of Assumption˜3 hold, and $\tau_{\varphi,d}=\sigma_{d}^{-1}$ by ˜1, we may apply Theorem˜1, which yields the desired result. ∎
C.3 Proofs of Section˜4
@@
$\Rightarrow$ Assume $p(x,z)=\varphi(x)^{\mathsf{\scriptscriptstyle T}}\psi(z)$. Then, by Eq.˜11, for all $h\in L_{2}(X)$,
@@
In order to prove Assumption˜5 we prove intermediate results.
@@
We first show that $\sigma_{d}^{-1}\leq\tau_{\hat{\varphi},d}\leq(\sigma_{d}-2\hat{\varepsilon}_{d})^{-1}$. First note that if $h\in\mathcal{H}_{\hat{\varphi},d}$, then by Proposition 6, $h\in\mathcal{V}_{d}$. Using the SVD of $\hat{\mathcal{T}}_{d}$ from Eq.˜15, and writing $h=\sum_{i=1}^{d}\langle h,\hat{v}_{i}\rangle\hat{v}_{i}$, we obtain:
@@
We now prove the equivalence between Eq.˜6 and Eq.˜7. While the proof strategy follows that of Kostic et al. [25, Theorem 1], we include our own version for completeness, as the parametrization of the learned operator differs. Specifically, [25] directly approximate the truncated operator $\mathcal{T}_{d}$ using a singular value decomposition of the form $\mathbbm{1}_{\mathcal{Z}}\otimes\mathbbm{1}_{\mathcal{X}}+\sum_{i=2}^{d}\hat{\sigma}_{i}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$, where the singular values $\hat{\sigma}_{i}$ are learned explicitly via a separate network. In contrast, our approach learns only the feature maps $(\varphi_{i},\psi_{i})$ and defines the approximation $\mathcal{T}_{d}(\hat{\varphi},\hat{\psi})=\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\varphi}_{i}$. As such, we provide a short self-contained derivation adapted to our setting for clarity. Recall that the Eckart–Young–Mirsky formulation of the objective is defined as
@@
Appendix D Extended Related Work
We now discuss the various ideas that have been proposed to solve NPIV problems. Extensive benchmarking for these methods has already been conducted in prior work; refer, for instance, to [36].
@@ Appendix D Extended Related Work
Building on similar principles, [18] proposed nonparametric methods based on both kernel555”Kernel” is meant here in the sense of density estimation, and not reproducing kernel: see [18, eq. (2.4)]. techniques and orthogonal series for estimating regression functions with instrumental variables. For their orthogonal series estimator, they transformed the marginal distributions of $X$ and $Z$ to be uniform and used Fourier expansions. The estimated Fourier coefficients of the structural function were obtained via a regularized regression involving estimated coefficients from the first stage. They were among the first to derive optimal convergence rates for this class of problems, explicitly linking these rates to the “difficulty” of the ill-posed inverse problem, which is characterized by the eigenvalues of the underlying integral operator.
@@
[35] introduced Kernel Instrumental Variable Regression (KIV), a direct nonparametric generalization of 2SLS. KIV models the relationships between the different variables as nonlinear functions in reproducing kernel Hilbert spaces (RKHSs). In stage 1, it learns a conditional mean embedding $\mu(z)=\mathbb{E}\left[\varphi(X)\mathrel{}\middle|\mathrel{}Z=z\right]$ where $\varphi(X)$ represents features of $X$ mapped into an RKHS $\mathcal{H}_{X}$. This learning is framed as a vector-valued kernel ridge regression, effectively estimating the conditional expectation operator $E:\mathcal{H}_{X}\to\mathcal{H}_{Z}$. In stage 2, a scalar-valued kernel ridge regression of the outcome $Y$ on the estimated means $\hat{\mu}(Z)$ is performed to estimate the structural function. The authors prove the consistency of KIV in the projected norm under mild conditions and derive conditions under which KIV achieves minimax optimal convergence rates. The analysis was later improved by [31], who established convergence in $L_{2}$ norm rather than the projected norm.
Another related approach, “fast IV” (FIV), was proposed in [39]. FIV studies nonlinear IV with high-dimensional instruments and proposes a two-stage pipeline that keeps the outcome model in a fixed RKHS/GP space but learns the first-stage kernel from a black-box adaptive regressor (e.g., a neural network) distilled into a compact kernel basis for estimating the conditional expectation operator. The resulting estimator plugs this learned kernel into standard kernelized IV schemes, achieving rates that adapt to the dimensionality of informative instrument features. Their work focuses on learning a kernel basis for the instrument. In contrast, our method learns paired spectral features for both the instrument and the treatment to approximate the top singular subspaces of the conditional expectation operator.
Saddle-point approaches.
@@
About the contrastive loss.
@@
Appendix E Experimental Details
In this section, we provide additional details for the experiments presented in the main text. Let $(Z,X)$ be generated as described in Section˜6 with $n=100000$. Figure˜4 displays both the true data-generating density and the density corresponding to a set of learned spectral features. Even for $d=11$, the resulting densities are complex. While it is feasible to conduct analogous experiments with higher values of $d$, we observed no qualitative changes in the outcomes. However, training and hyperparameter tuning became increasingly challenging as $d$ grew.
@@ E.1 Models Employed
E.2 Expanded example of the ugly scenario
@@
mirroring the setup from Section˜6. We vary the spectrum by setting $\sigma_{1:k}=c$ and $\sigma_{k+1:d-1}=0$ for $k\in\{1,\ldots,d-1\}$. We then define the target function:
@@
which is uniformly spread across the feature directions. As $k$ increases, the projection of $h_{0}$ onto the singular space of $\mathcal{T}$ increases in discrete steps from 0 to 1. This results in a corresponding qualitative improvement in the accuracy of the 2SLS estimator for $h_{0}$. Figure˜5 illustrates this behavior: for small values of $k$, where $h_{0}$ is largely orthogonal to the singular functions of $\mathcal{T}$, the estimate fails to recover $h_{0}$. As $k$ grows, and more of $h_{0}$’s energy lies in the span of these singular functions, the 2SLS estimate increasingly aligns with the true function.
@@
Appendix F Estimating spectral alignment
In this section, we present the details of the methodology introduced in Section˜5. We recall the central formula provided in Eq.˜10.
$\langle v_{i},h_{0}\rangle_{L_{2}(X)}=\frac{1}{\sigma_{i}}\mathbb{E}[Y\cdot u_{i}(Z)].$ (16)
To evaluate spectral alignment, we shall approximate the RHS of the above. Suppose we have learned an operator $\widehat{\mathcal{T}}_{d}=\sum_{i=1}^{d}\hat{\psi}_{i}\otimes\hat{\phi}_{i}=\Psi\Phi^{*}$ which approximates the true conditional mean operator $\mathcal{T}$. Here $\Psi\colon\mathbb{R}^{d}\to L^{2}(Z)$ and $\Phi\colon\mathbb{R}^{d}\to L^{2}(X)$ are the maps sending $\alpha\stackrel{{\scriptstyle\Psi}}{{\mapsto}}\sum_{i=1}^{d}\alpha_{i}\hat{\psi}_{i}$ and $\alpha\stackrel{{\scriptstyle\Phi}}{{\mapsto}}\sum_{i=1}^{d}\alpha_{i}\hat{\phi}_{i}$. One can compute the SVD of $\widehat{\mathcal{T}}_{d}$ in two steps. We start with the derivations assuming access to population quantities. Let $C_{Z}$, and $C_{X}$ be the $d\times d$ population covariance matrices of the learned $Z$, and $X$ features, respectively. For any $z$ and $x$, we then introduce the whitened feature vectors $\psi_{i}^{\prime}(z)=(C_{Z}^{-1/2}\hat{\psi}(z))_{i}$ and $\phi_{i}^{\prime}(x)=(C_{X}^{-1/2}\hat{\phi}(x))_{i}$. The corresponding operators, $\Psi^{\prime}=\Psi C_{Z}^{-1/2}$ and $\Phi^{\prime}=\Phi C_{X}^{-1/2}$, are isometries. We can then write